Nnesha (nnesha):

need help pleassee help me A school is ordering boxes of copy paper for its teachers. When placing their order, the paper company explains the pricing to them as follows. If they order no more than 1000 boxes, the company will charge them $35 per box. If they order more than 1000 boxes, the company will charge them $27.50 per box. Which type of function would best model this situation: a step function or a piecewise function that is not a step function? Explain.

OpenStudy (broskishelleh):

This is a step function since it has a finite number of pieces, namely f(x) = 35*x if x < or = 1000 = 27.50*x if x > 1000

OpenStudy (broskishelleh):

@Nn

OpenStudy (broskishelleh):

@Nnesha

Nnesha (nnesha):

ok thanks

OpenStudy (broskishelleh):

Can I have a medal and fan

OpenStudy (broskishelleh):

f(x)={50,50+0.1(x−500),if x≤500if x>500a

OpenStudy (broskishelleh):

we need a piecewise function that is not a greatest integer function because the phone bill can be in cents also... for example, if u have used 501 minutes, f(x) will be 50+0.1(501-500) = 50.1. which is clearly not an integer. so you need a piecewise function which is not a greatest integer funciton

OpenStudy (broskishelleh):

@Nnesha

OpenStudy (broskishelleh):

@Nnesha Do you understand?

Nnesha (nnesha):

i already have that answer exactly same like that

OpenStudy (broskishelleh):



OpenStudy (broskishelleh):

Well that was for almost no reason at all

Nnesha (nnesha):

what is that

Nnesha (nnesha):

what

OpenStudy (broskishelleh):

Can I have a fan? @Nnesha

OpenStudy (broskishelleh):

@Nnesha If you fan me, you can tag me into your questions

Nnesha (nnesha):

now happy ........i just start fanning people

Nnesha (nnesha):

LOL